OpenStudy (anonymous):

A gold nugget has a volume of 20.0 cm3 and its density is 19.3 g/cm3. What is the mass of the gold nugget?

OpenStudy (anonymous):

use the density = mass / volume

OpenStudy (anonymous):

\[D = \frac{ mass }{ volume } \] the density for the gold nugget \[\frac{ mass }{ volume } =\frac{ 19.3g }{ cm ^{3} } \] you are given the volume for the gold nugget but the mass is unknown. The following equation can be written: \[\frac{ 19.3g }{ 1 cm ^{3} } = \frac{ m }{ 20.0cm ^{3} }\] multiply both sides by 20.0cm^3 (on the right side 20cm^3/20cm^3 will cancel each other out and you'll be left eith only m. On the left side the the cm^3 in the numerator and denominator will cancel each other out, and you'll be left with grams) The equation will be 20.0 * 19.3 = m